3.) The Electric Parts Division of Babylon Corporation has recently made a proposal to Cable division. The Electric Parts Division would like to make a ‘special switch’ with gold-plated parts. It would like the Cable Division to make the 4000 units of gold-plated parts and provide them to the Electric Parts Division at a transfer price of $60 per unit. The estimated variable cost per unit would be $44.

Cable Division’s standard unit sells for $30 per unit and has variable costs of $17.

  1. Compute the minimum transfer price that the Electric Parts Division should be willing to accept, if the Cable Division has excess capacity.
  2. Compute the minimum transfer price that the Cable Division should be willing to accept, if it is currently operating at full capacity and if it has to forgo the sales of 5,000 units of standard unit.
  3. In the situation explained in part b, if the Cable Division doesn’t accept the offer of Electric Parts Division, and Electric Parts division purchases the parts from an outside supplier at a cost of $65 per unit, what will be the results of that for the overall company?
Answer #1
Req a.
Minimum transfer price in case of excess capacity exist is variable cost of manufacturing.
Therefore,
Minimum transfer price is $ 44 perunit
Req b.
However, when no excess capacity exist, the transfer price shall be variable cost and loss of contribution:
Contribution per unit = (30-17) 13
Multiply: Loss of Sales units 5000
Total loss of contribution 65000
Divide: Units to be transferred 4000
Loss of contribution per unit 16.25
Min transfer price = 16.25+44 = 60.25
Req c.
Loss per unit to the company = 65-60.25 = 4.75 per unit
Total loss to the company = 4.75*4000 = $ 19000